40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 7 17:39:40 2012, keiko, Update, LSC, MICH noise budget on 5 Mar Mar5-MICHbudget.pngMar5-MICHbudget.png
    Reply  Wed Mar 7 20:53:13 2012, keiko, Update, LSC, MICH noise budget on 5 Mar Mar6-MICHbudget.png
    Reply  Thu Mar 8 00:57:48 2012, keiko, Update, LSC, MICH noise budget on Mar 5, Mar 6, and old Mar7MICHbudgettotal.png
Message ID: 6380     Entry time: Wed Mar 7 20:53:13 2012     In reply to: 6376
Author: keiko 
Type: Update 
Category: LSC 
Subject: MICH noise budget on 5 Mar 

 

 Mar6-MICHbudget.png

This is the MICH noise budget on 6th March. 1Hz peak got a bit better as the BS sus control gain was increased.

 

ELOG V3.1.3-